OpenStudy (anonymous):

The coordinates below represent two linear equations. How many solutions does this system of equations have? Line 1 x y –2 7 0 6 Line 2 x y 3 0 0 –3 A. 0 B. exactly 1 C. exactly 2 D. infinitely many For line 1 I got -1/2 For line 2 I got -1

OpenStudy (anonymous):

@ispike

OpenStudy (anonymous):

sorry i would have to look back into this its been a long time since of done linear equations

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@AnswerMyQuestions

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

\[y-y _{1}=\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\left( x-x _{1} \right)\]

OpenStudy (anonymous):

i got B